OpenStudy (anonymous):

Find the value of x. Round your answer to the nearest tenth. Tan46 = X over 12

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \tan(46) = \frac{x}{12}\)

ganeshie8 (ganeshie8):

to find x, we need to isolate x. so wat should be our first step ?

ganeshie8 (ganeshie8):

multiply 12 both sides ?

OpenStudy (anonymous):

yes multiply.

ganeshie8 (ganeshie8):

\(\large \large 12 \times \tan(46) = 12 \times \frac{x}{12}\)

OpenStudy (anonymous):

so 12 gets crossed out right.

ganeshie8 (ganeshie8):

\(\large \large 12 \times \tan(46) = \cancel{12} \times \frac{x}{\cancel{12}}\)

ganeshie8 (ganeshie8):

yes crossed out

ganeshie8 (ganeshie8):

\(\large \large 12 \times \tan(46) = x\)

ganeshie8 (ganeshie8):

\(\large x = \large 12 \times \tan(46) \)

OpenStudy (anonymous):

im not sure what to do in that step

ganeshie8 (ganeshie8):

now, use ur calculator and find out wat tan(46) is

ganeshie8 (ganeshie8):

then multiply it wid 12

ganeshie8 (ganeshie8):

u knw how to find tan(46) in ur calculator ?

OpenStudy (anonymous):

tan of 46 = 1.03 right?

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

finish it off,

OpenStudy (anonymous):

12.3?

ganeshie8 (ganeshie8):

\(\large x = 12 \times \tan(46)\) \(\large = 12 \times 1.03\) \(\large = 12.36\) \(\large \sim 12.4\)

ganeshie8 (ganeshie8):

when u round 12.36, it becomes 12.4

OpenStudy (anonymous):

thank you! :)

ganeshie8 (ganeshie8):

np you're wlcme :)